# Quantile Activation: Correcting a failure mode of ML models

### v2

@@ Appendix C Extended Results Section
Figure 7: Comparing QAct with ReLU activation and DINOv2 (small).
@@
Figure 7 shows the corresponding results. The first experiment compares QAct + watershed with ReLU + Cross-Entropy on two standard networks – ResNet18 and DenseNet. With respect to accuracy, we observe that while at severity 00, ReLU + Cross-Entropy slightly outperforms QAct + watershed, as severity increases QAct + watershed is far more stable. We even outperform DINOv2(small) (22M parameters) at severity $5$. Moreover, with respect to calibration error, we see a consistent trend across distortions. As (Challa et al., 2023) argues, this helps in building more robust systems compared to one where calibration error increases across distortions.
@@
Figure 8: Triplet vs Watershed vs Cross-Entropy
@@
Figure 9: Watershed vs Cross-Entropy when using ReLU activation
@@
Figure 11: QActvs ReLU vs pReLU vs Selu activations on ResNet18
@@
Figure 12: QActvs ReLU vs pReLU vs Selu activations on Densenet
@@
Figure 13: QActvs ReLU on Resnet18+CIFAR100
@@
Figure 14: Effect of Quantile Classifier. We use ResNet18 and DinoV2 architectures on CIFAR10.
@@
Figure 15: Effect of Quantile Classifier. We use ResNet18 and DinoV2 architectures on CIFAR100.
@@
To measure the robustness of the proposed method we use the metric Acc@Dist_i - Acc@Dist_j which measures the drop in accuracy when the distortion severity is increased from $i\to j$. A method is considered to be better if the values are lower, i.e if the drop in accuracy is smaller than comparitive methods. Tables 1 and 2 shows the comparison between ReLU, QAct and DinoV2(s). We see that, in all the cases QAct outperforms both ReLU and DinoV2(s) at all possible $i\to j$.
@@ Appendix D Watershed Loss
The authors in (Challa et al., 2024) proposed a novel classifier – watershed classifier, which works by learning similarities instead of the boundaries. Below we give the brief idea of the loss function, and refer the reader to the original paper for further details.
@@
Observe that the watershed loss is very similar to metric learning losses. The authors in (Challa et al., 2024) claim that this offers better generalization, and show that this is consistent with 1NN classifier. Moreover, they show that this classifier (without considering $f_{\theta}$) has a VC dimension which is equal to the number of classes. While metric learning losses are similar, there is no such guarantee with respect to classification. This motivated our choice of using watershed loss over other metric learning losses.

### v3

@@ 5 Conclusion And Future Work
Acknowledgments
Aditya Challa and Laurent Najman acknowledges the support from CEFIPRA(68T05-1) . Snehanshu Saha, Aditya Challa, Sravan Danda would like to thank the Anuradha and Prashanth Palakurthi Center for Artificial Intelligence Research (APPCAIR) and ANRF CRG-DST (CRG/2023/003210) for support. Snehanshu Saha acknowledges SERB SURE-DST (SUR/2022/001965) and the DBT-Builder project (BT/INF/22/SP42543/2021), Govt. of India for partial support. Sravan Danda would like to thank ANRF for funding from SERB-SURE project number SUR/2022/002735.
@@ Appendix C Extended Results Section
Figure 7: Comparing QAct with ReLU activation and DINOv2 (small) on CIFAR10C
@@
Figure 7 shows the corresponding results. The first experiment compares QAct + watershed with ReLU + Cross-Entropy on two standard networks – ResNet18 and DenseNet. With respect to accuracy, we observe that while at severity 00, ReLU + Cross-Entropy slightly outperforms QAct + watershed, as severity increases QAct + watershed is far more stable. We even outperform DINOv2(small) (22M parameters) at severity $5$. Moreover, with respect to calibration error, we see a consistent trend across distortions. As (Challa et al., 2024b) argues, this helps in building more robust systems compared to one where calibration error increases across distortions.
@@
Figure 8: Triplet vs Watershed vs Cross-Entropy using Resnet18+CIFAR10C
Remark:
@@
Figure 9: Watershed vs Cross-Entropy when using ReLU activation using CIFAR10C
@@
Figure 11: QActvs ReLU vs pReLU vs Selu activations on ResNet18 + CIFAR10C
@@
Figure 12: QActvs ReLU vs pReLU vs Selu activations on Densenet+CIFAR10C
@@
Figure 13: QActvs ReLU on Resnet18+CIFAR100C/TinyImagenetC
@@
Figure 14: Effect of Quantile Classifier. We use ResNet18 and DinoV2 architectures on CIFAR10C.
@@
Figure 15: Effect of Quantile Classifier. We use ResNet18 and DinoV2 architectures on CIFAR100C.
@@
To measure the robustness of the proposed method we use the metric Acc@Dist_i - Acc@Dist_j which measures the drop in accuracy when the distortion severity is increased from $i\to j$. A method is considered to be better if the values are lower, i.e if the drop in accuracy is smaller than comparitive methods. Tables 2 and 3 shows the comparison between ReLU, QAct and DinoV2(s). We see that, in all the cases QAct outperforms both ReLU and DinoV2(s) at all possible $i\to j$.
@@ Appendix D Watershed Loss
The authors in (Challa et al., 2024a) proposed a novel classifier – watershed classifier, which works by learning similarities instead of the boundaries. Below we give the brief idea of the loss function, and refer the reader to the original paper for further details.
@@
Observe that the watershed loss is very similar to metric learning losses. The authors in (Challa et al., 2024a) claim that this offers better generalization, and show that this is consistent with 1NN classifier. Moreover, they show that this classifier (without considering $f_{\theta}$) has a VC dimension which is equal to the number of classes. While metric learning losses are similar, there is no such guarantee with respect to classification. This motivated our choice of using watershed loss over other metric learning losses.
Appendix E Formal Explanation of Failure Mode in Linear Models and why Quantiles fix it
Here, we discuss the failure mode presented in section 1 formally in the context of linear models. And then show, why the quantile activation as presented corrects this.
Linear Models under simple Distribution Shift:
Let $\{({\bm{x}}_{i},y_{i})\}$ denote the set of examples from the joint probability distribution of $p({\textnormal{x}},{\textnormal{y}})$. Assume for simplicity the binary classification problem i.e $y_{i}\in\{0,1\}$. Let $w^{t}{\bm{x}}+b$ denote the linear classifier trained on this. Note that the classification rule is then simply $I[w^{t}{\bm{x}}+b\geq 0]$.
Now let ${\bm{A}}$ denote a transformation matrix. Let ${\mathbf{x}}$ (marginal) be transformed as ${\bm{A}}{\mathbf{x}}$. If ${\bm{x}}^{\prime}={\bm{A}}{\bm{x}}$ denote a sample from the transformed distribution, the classification rule applied to this sample is
$I[w^{t}{\bm{x}}^{\prime}+b\geq 0]=I[w^{t}({\bm{A}}{\bm{x}})+b\geq 0]$ (13)
It is easy to see that, even under the simple case of ${\bm{A}}=c{\bm{I}}$, the class assignment can vary arbitrarily with $c$.
Linear Models + Quantile Activation under simple Distribution Shift:
What we propose in the article, is to instead consider the following classification rule. If $x^{\prime}={\bm{A}}x$ denote a sample from the transformed distribution,
$P(w^{t}{\bm{x}}^{\prime}>w^{t}({\bm{A}}{\mathbf{x}}))\geq 0.5=P(w^{t}({\bm{A}}{\bm{x}})>w^{t}({\bm{A}}{\mathbf{x}}))\geq 0.5$ (14)
Here, if ${\bm{A}}=c{\bm{I}}$, then it is easy to see that the class assignment is independent of the value of c. More formally, we have the following proposition.
Proposition 1.
Let $A$ denote the transformation as described above. Then, as long as ${\bm{A}}^{t}w=\alpha w$ with $\alpha>0$ then the class assignment of the linear model with quantile activation does not change under the transformation ${\bm{A}}{\bm{x}}+z$ where $z$ can be arbitrary.
The proof follows by simple substitution in equation 14.